OpenStudy (anonymous):

I have problem with bracket: My proble is this: for example I have: 2x-2 and 5x-5 when I must supract 2x-2 with 5x-5 I must write them so: 2x-2-(5x-5) or so: 2x-2-5x-5 ? or this example: 2x with 5x-5 2x-(5x-5) or 2x-5x-5 thx

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[2x-2-(5x-5)\] is the question?

OpenStudy (anonymous):

ehm no I have writen the question....

OpenStudy (misty1212):

oh

OpenStudy (mathmath333):

u should write it in brackets as this \(2x-2-(5x-5)\) if bracket is simplied it changes to this \(2x-2-5x+5)\)

OpenStudy (misty1212):

i have to say "when I must supract 2x-2 with 5x-5 " doesn't mean much you can subtract \(5x-5\) FROM \(2x-2\) or you can subtract \(2x-2\) FROM \(5x-5\)

OpenStudy (anonymous):

a ok thanks

OpenStudy (misty1212):

if it is subtract \(5x-5\) FROM \(2x-2\) then \[2x-2-(5x-5)\] or \[2x-2-5x+5\]

OpenStudy (anonymous):

ok thanks I understand

OpenStudy (mathmath333):

remember the rule \(\large \begin{align} \color{black}{a-(b-c)=a-b+c\neq a-b-c \hspace{.33em}\\~\\}\end{align}\)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Ok thx I go by :)

OpenStudy (mathmath333):

\(\Huge\tt \begin{align} \color{red}{\ddot \smile \hspace{.33em}\\~\\}\end{align}\)